# The formation of a soliton gas condensate for the focusing Nonlinear Schrödinger equation

### v1

@@ 6.1 Local parametrix near $z={\overline{A}}$
$E_{{\overline{A}}}(z)=Q_{{\overline{A}}}(z)e^{i\pi\sigma_{3}/4}\,\Psi\left(\varphi_{{\overline{A}}}(z)\right)\,e^{-i\pi\sigma_{3}/4}e^{\ln(N)\frac{2c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}\sigma_{3}}\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{-\sigma_{3}/2}$ (6.3)
@@
$\varphi_{{\overline{A}}}(z)^{1/2}=-\ln(N)\frac{c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}\,.$ (6.4)
@@
Notice that we can factorize the matrix $\mathfrak{N}^{+}(z)$ as follows
$\mathfrak{N}^{+}(z)=\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}\begin{pmatrix}1&0\\ -ie^{\ln(N)\tilde{u}(z)}&1\end{pmatrix}\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{-\sigma_{3}/2}$
and therefore, using conditions (4.13) and the definition (4.6), we can rewrite the jump of $E_{{\overline{A}}}$ as
$J_{E_{{\overline{A}}}}(z)=\Big{(}e^{2\widetilde{y}_{-}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}\left\{\begin{array}[]{lr}\begin{pmatrix}0&i\\ i&0\end{pmatrix},&z\in\eta_{4}\\ \\[4.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\frac{4c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}}&1\end{pmatrix},&z\in\omega_{1}\cup\omega_{2}\end{array}\right\}\Big{(}e^{2\widetilde{y}_{+}(z)}\mathcal{Y}(z)\Big{)}^{-\sigma_{3}/2}.$ (6.8)
Next, we consider the function
$P(z)=E_{{\overline{A}}}(z)\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}$ (6.9)
and it is easy to check directly from the jumps of $E_{{\overline{A}}}$, that $P$ satisfies the following jump relation
$P_{+}(z)=P_{-}(z)\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},&z\in\eta_{4}\\ \\[4.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\frac{4c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}}&1\end{pmatrix},&z\in\omega_{1}\cup\omega_{2}\end{cases}$ (6.10)
Analysis of the local behavior of the function $\frac{4c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}$ near $z={\overline{A}}$ , prompts to introduce locally the map $\varphi_{{\overline{A}}}:z\mapsto\varphi$ defined via the relation
$\left(\varphi_{{\overline{A}}}(z)\right)^{1/2}\equiv\varphi^{1/2}=-\ln(N)\frac{c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}\,.$ (6.11)
The jumps of $P$ then become
$J_{P}(\varphi_{{\overline{A}}}(z))=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&(-\infty,0)\\ \\[4.0pt] \begin{pmatrix}1&0\\ -ie^{-4\varphi^{1/2}}&1\end{pmatrix},\quad&\varphi_{{\overline{A}}}(\omega_{1})\cup\varphi_{{\overline{A}}}(\omega_{2})\end{cases}$
where $\varphi_{{\overline{A}}}({\omega}_{1})$ and $\varphi_{{\overline{A}}}({\omega}_{2})$ are the images of $\omega_{1}$ and $\omega_{2}$ under the map $\varphi_{{\overline{A}}}$. We then define the function
$\Psi_{{\overline{A}}}(\varphi_{{\overline{A}}}(z))=P(\varphi_{{\overline{A}}}(z))\,e^{2\varphi^{1/2}\,\sigma_{3}}$ (6.12)
which satisfies the jump relation
$\Psi_{{\overline{A}}+}(\varphi_{{\overline{A}}}(z))=\Psi_{{\overline{A}}-}(\varphi_{{\overline{A}}}(z))\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&(-\infty,0)\\ \\[4.0pt] \begin{pmatrix}1&0\\ -i&1\end{pmatrix},\quad&\varphi_{{\overline{A}}}(\omega_{1})\cup\varphi_{{\overline{A}}}(\omega_{2})\end{cases}$ (6.13)
Finally, consider the function
$\Psi_{{\overline{A}}}(\varphi_{{\overline{A}}}(z))=e^{i\pi\sigma_{3}/4}\,\Psi(\varphi_{{\overline{A}}}(z))\,e^{-i\pi\sigma_{3}/4}$ (6.14)
where $\Psi$ is given by equation (6.2). We can check directly by using the jump relation of $\Psi$ that $\Psi_{{\overline{A}}}$ in (6.14) satisfies the desired jumps. Unraveling the transformations, we see that (6.3) satisfies the desired jump relations near $z={\overline{A}}$, for any analytic and invertible matrix valued function $Q_{{\overline{A}}}$. The quantity $Q_{{\overline{A}}}$ will be determined at a later stage. ∎
6.2 Local parametrix near $z=-A$
Recall that $\mathcal{U}_{-A}$ denotes a small neighborhood of $z=-A$ and $E_{-A}$ is the solution to the local problem near $z=-A$. Inside $\mathcal{U}_{-A}$, $E_{-A}$ satisfies the jump relation
$E_{-A+}(z)=E_{-A-}(z)J_{E_{-A}}(z),\quad J_{E_{-A}}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},&\quad z\in\eta_{4}\\ \\[2.0pt] e^{\left(-\Delta+\frac{2}{\tau}\ln(N)\right)\sigma_{3}},&\quad z\in\gamma\\ \\[2.0pt] \mathfrak{N}^{+}(z),&\quad z\in\omega_{1}\cup\omega_{2}\end{cases}$ (6.15)
Proposition 6.2.
The matrix valued function $E_{-A}$ given by the expression
$E_{-A}(z)=Q_{-A}(z)\,e^{i\pi\sigma_{3}/4}\,\Psi(\varphi_{-A}(z))\,e^{-i\pi\sigma_{3}/4}e^{\ln(N)\frac{2c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{R(s)}\sigma_{3}}\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{\sigma_{3}/2}\begin{cases}e^{-(\widetilde{y}(z)-\frac{1}{\tau}\ln(N))\sigma_{3}},\quad z\in\Omega\\ e^{-(\widetilde{y}(z)+\frac{1}{\tau}\ln(N))\sigma_{3}},\quad z\in\widetilde{\Omega}\end{cases}$ (6.16)
satisfies the jump relation (6.15). Here, $\Psi$ is as in (6.2), $Q_{-A}$ is an analytic function inside $\mathcal{U}_{-A}$, the map $\varphi_{-A}$ is defined as
$\varphi_{-A}(z)^{1/2}=-\ln(N)\frac{c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{R(s)}\,,$ (6.17)
and the domains $\Omega,\,{\widetilde{\Omega}}$ are as in Figure 10.
Proof.
We introduce the function $m$:
$m(z)=\begin{cases}E_{-A}(z)\,e^{\left(\widetilde{y}(z)-\frac{1}{\tau}\ln(N)\right)\sigma_{3}},\quad&z\in\Omega\\ \\[2.0pt] E_{-A}(z)\,e^{\left(\widetilde{y}(z)+\frac{1}{\tau}\ln(N)\right)\sigma_{3}},\quad&z\in\widetilde{\Omega}\end{cases}\,.$ (6.18)
$\gamma$$\eta_{4}$$\omega_{2}$$\omega_{1}$$\Omega$$\tilde{\Omega}$
Figure 10: Contour for RHP 6.15
$m_{+}(z)=m_{-}(z)J_{m}(z),$ (6.19)
and the jump matrix $J_{m}(z)$ can be factored as follows
$J_{m}(z)=\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{-\sigma_{3}/2}\left\{\begin{array}[]{lr}\begin{pmatrix}0&i\\ i&0\end{pmatrix},&z\in\eta_{4}\\ \\[2.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\frac{4c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{R(s)}}&1\end{pmatrix},&z\in\omega_{1}\cup\omega_{2}\end{array}\right\}\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{\sigma_{3}/2}\,,$ (6.23)
and we define the function
$P(z)=m(z)\,\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{-\sigma_{3}/2}\,,$ (6.24)
which satisfies the jump relations
$P_{+}(z)=P_{-}(z)\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{4}\\ \\[2.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\frac{4c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{R(s)}}&1\end{pmatrix},\quad&z\in\omega_{1}\cup\omega_{2}\end{cases}\,.$
Analysis of the local behavior of the function $\frac{4c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{R(s)}$ near $z=-A$ , prompts to introduce locally the map $\varphi_{-A}:z\mapsto\varphi$ defined via the relation
$\left(\varphi_{-A}(z)\right)^{1/2}\equiv\varphi^{1/2}=-\ln(N)\frac{c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{R(s)}\,,$ (6.25)
and therefore the jumps of $P$ become
$J_{P}(\varphi_{-A}(z))=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&(-\infty,0)\\ \\[4.0pt] \begin{pmatrix}1&0\\ -ie^{-4\varphi^{1/2}}&1\end{pmatrix},\quad&\varphi_{-A}(\omega_{1})\cup\varphi_{-A}(\omega_{2}).\end{cases}$
We then define the function
$\Psi_{-A}(\varphi_{-A}(z))=P(\varphi_{-A}(z))\,e^{2\varphi^{1/2}\,\sigma_{3}}\,,$ (6.26)
which satisfies the jump relation
$\Psi_{-A+}(\varphi_{-A}(z))=\Psi_{-A-}(\varphi_{-A}(z))\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&(-\infty,0)\\ \\[4.0pt] \begin{pmatrix}1&0\\ -i&1\end{pmatrix},\quad&\varphi_{-A}(\omega_{1})\cup\varphi_{-A}(\omega_{2})\end{cases}\,.$ (6.27)
Notice that this is the jump relation that $\Psi_{{\overline{A}}}$ satisfies. Therefore, the function $\Psi_{-A}$ is given by
$\Psi_{-A}(\varphi_{-A}(z))=e^{i\pi\sigma_{3}/4}\,\Psi(\varphi_{-A}(z))\,e^{-i\pi\sigma_{3}/4}\,,$ (6.28)
and we can check directly by using the jumps of $\Psi$, that $\Psi_{-A}$ above, satisfies the desired jumps. Tracking the subsequent transformations we have completed the proof of the proposition. ∎
@@ 6.3 Local parametrix near $z=A$
Recall that $\mathcal{U}_{A}$ is a small neighborhood of $z=A$ and we denote by $E_{A}$ the solution to the local problem near $z=A$. Inside $\mathcal{U}_{A}$, $E_{A}$ satisfies the jump relation
$E_{A+}(z)=E_{A+}(z)J_{E_{A}}(z),\quad J_{E_{A}}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{3}\\ \\[2.0pt] \mathfrak{M}^{-}(z),\quad&z\in\omega_{3}\cup\omega_{4}\end{cases}\,.$ (6.29)
@@
$E_{A}(z)=Q_{A}(z)\Psi_{A}\left(\varphi_{A}(z)\right)e^{\ln(N)\frac{2c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{R(s)}\sigma_{3}}\Big{(}\frac{e^{2\widetilde{y}(z)}}{\mathcal{Y}(z)e^{4i\theta(z)}}\Big{)}^{-\sigma_{3}/2}\,,$ (6.30)
satisfies the jump relation (6.29). Here, the function $\Psi_{A}$ is given by the expression
$\Psi_{A}(\varphi_{A}(z))=-\sigma_{2}\,\Psi_{{\overline{A}}}(\varphi_{A}(z))\,\sigma_{2}\,,\quad\sigma_{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\,,$ (6.31)
and satisfies the jumps shown in Fig. 11. Moreover, $\Psi_{{\overline{A}}}$ is as in (6.14) and satisfies the jump relation (6.13), $Q_{A}$ is an analytic function inside $\mathcal{U}_{A}$, and the map $\varphi_{A}$ is defined as
$\varphi_{A}(z)^{1/2}=-\ln(N)\frac{c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{R(s)}\,.$ (6.32)
@@
Similar analysis as before, simplifies $J_{E_{A}}$ into
$J_{E_{A}}(z)=\Bigg{(}\frac{e^{2\widetilde{y}_{-}(z)}}{\mathcal{Y}(z)e^{4i\theta(z)}}\Bigg{)}^{\sigma_{3}/2}\left\{\begin{array}[]{lr}\begin{pmatrix}0&i\\ i&0\end{pmatrix},&z\in\eta_{3}\\ \\[4.0pt] \begin{pmatrix}1&-ie^{-\ln(N)\frac{4c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{R(s)}}\\ 0&1\end{pmatrix},&z\in\omega_{3}\cup\omega_{4}\end{array}\right\}\Bigg{(}\frac{e^{2\widetilde{y}_{+}(z)}}{\mathcal{Y}(z)e^{4i\theta(z)}}\Bigg{)}^{-\sigma_{3}/2}\,.$ (6.36)
We define the matrix function
$P(z)=E_{A}(z)\,\Bigg{(}\frac{e^{2\widetilde{y}(z)}}{\mathcal{Y}(z)e^{4i\theta(z)}}\Bigg{)}^{\sigma_{3}/2}\,,$ (6.37)
and we introduce locally the map $\varphi_{A}:z\mapsto\varphi$ via the relation
$\left(\varphi_{A}(z)\right)^{1/2}\equiv\varphi^{1/2}=-\ln(N)\frac{c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{R(s)}\,.$ (6.38)
The jumps of $P$ now become
$J_{P}(\varphi_{A}(z))=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&(0,-\infty)\\ \\[4.0pt] \begin{pmatrix}1&-ie^{4\varphi^{1/2}}\\ 0&1\end{pmatrix},\quad&\varphi_{A}(\omega_{3})\cup\varphi_{A}(\omega_{4})\end{cases}\,.$ (6.39)
Finally, we define the function
$\Psi_{A}(\varphi_{A}(z))=P(\varphi_{A}(z))\,e^{2\varphi^{1/2}\,\sigma_{3}}\,,$ (6.40)
which satisfies the jumps shown in Fig. 11 with the prescribed orientation. We claim that the function $\Psi_{A}$ given by the expression
$\Psi_{A}(\varphi_{A}(z))=-\sigma_{2}\,\Psi_{{\overline{A}}}(\varphi_{A}(z))\,\sigma_{2}$ (6.41)
satisfies the jumps shown in Fig. 11, where $\Psi_{{\overline{A}}}$ is as in (6.14) and satisfies the jump relation (6.13).
[…]
Figure 11: Jumps for $\Psi_{A}$ on $\widetilde{\Sigma}$
To show this claim, let us denote by $\Sigma$ the contour of non-analyticity for $\Psi_{{\overline{A}}}$, and $\widetilde{\Sigma}$ the contour of non-analyticity for $\Psi_{A}$. Note that the two contours are oriented the opposite way.
[…]
Figure 12: Jumps for $\Psi_{{\overline{A}}}$ on $\Sigma$
It is easy to check that
$\Psi_{{\overline{A}}\pm}(z)\Big{|}_{\Sigma}=\Psi_{{\overline{A}}\mp}(z)\Big{|}_{\widetilde{\Sigma}}$
and therefore it is immediate that
$J_{\Psi_{{\overline{A}}}}(z)\Big{|}_{\widetilde{\Sigma}}=\left(J_{\Psi_{{\overline{A}}}}(z)\Big{|}_{\Sigma}\right)^{-1}$
which implies that $\Psi_{{\overline{A}}}$ satisfies the jumps shown on Fig. 13 along the contour $\widetilde{\Sigma}$.
[…]
Figure 13: Jumps for $\Psi_{{\overline{A}}}$ on $\widetilde{\Sigma}$
Using the jumps of $\Psi_{{\overline{A}}}$ on $\widetilde{\Sigma}$, it is easy to check that the function defined in (6.31) satisfies the desired jump relation on $\widetilde{\Sigma}$. Tracking the subsequent transformations we obtain the local solution given in (6.30). ∎
Similarly, one can construct the local solution of the problem near $z=-{\overline{A}}$, the details of which we omit.

### v2

@@ 6.2 Local parametrix near $z={\overline{A}}$
$E_{{\overline{A}}}(z)=Q_{{\overline{A}}}(z)e^{i\pi\sigma_{3}/4}\,\Psi\left(\varphi_{{\overline{A}}}(z)\right)\,e^{-i\pi\sigma_{3}/4}e^{-2\varphi_{{\overline{A}}}(z)^{1/2}\sigma_{3}}\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{-\sigma_{3}/2}$ (6.18)
@@
$\varphi_{{\overline{A}}}(z)^{1/2}=-\ln(N)\frac{c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{S(s)}\,.$ (6.19)
@@
It suffices to show that $E_{{\overline{A}}}(z)$ as in equation (6.18) satisfies the jump relation (6.17). We therefore compute the jump of $E_{{\overline{A}}}(z)$, first along $\eta_{4}$. Note that the non-analytic functions in the definition of $E_{{\overline{A}}}(z)$ along $\eta_{4}$ are: $\Psi\left(\varphi_{{\overline{A}}}(z)\right)$, $\varphi_{{\overline{A}}}(z)^{1/2}$ and $\widetilde{y}(z)$. We then have:
$J_{E_{{\overline{A}}}}(z)=\left[\left(E_{{\overline{A}}}\right)_{-}(z)\right]^{-1}\left(E_{{\overline{A}}}\right)_{+}(z)$
$=\left(e^{2\widetilde{y}_{-}(z)}\mathcal{Y}(z)\right)^{\sigma_{3}/2}e^{2\varphi_{{\overline{A}}-}(z)^{1/2}\sigma_{3}}e^{i\pi\sigma_{3}/4}J_{\Psi}\left(\varphi_{{\overline{A}}}(z)\right)e^{-i\pi\sigma_{3}/4}e^{-2\varphi_{{\overline{A}}+}(z)^{1/2}\sigma_{3}}\left(e^{2\widetilde{y}_{+}(z)}\mathcal{Y}(z)\right)^{\sigma_{3}/2}.$ (6.20)
From the properties of $\Psi$, for $z\in\eta_{4}$, we have:
$e^{i\pi\sigma_{3}/4}J_{\Psi}\left(\varphi_{{\overline{A}}}(z)\right)e^{-i\pi\sigma_{3}/4}=\begin{pmatrix}0&i\\ i&0\end{pmatrix}.$ (6.21)
In addition, $\varphi_{{\overline{A}}}(z)^{1/2}$ has jump across $\eta_{4}$, specifically:
$\varphi_{{\overline{A}}-}(z)^{1/2}=-\varphi_{{\overline{A}}+}(z)^{1/2},\quad z\in\eta_{4}\ .$
Finally, recall that the function $\widetilde{y}$ satisfies the jump relation (4.9) along $\eta_{4}$. Using all of these relations, straightforward multiplication of the right-hand side of equation (6.20) yields the desired jump of $E_{{\overline{A}}}(z)$ along $\eta_{4}$. Next, we compute the jump of $E_{{\overline{A}}}(z)$ along $\omega_{1}\cup\omega_{2}$. Note that the only non-analytic piece in the definition of $E_{{\overline{A}}}(z)$ along these two rays is the function $\Psi\left(\varphi_{{\overline{A}}}(z)\right)$. We then have:
$J_{E_{{\overline{A}}}}(z)$ $=\left(e^{2\widetilde{y}(z)}\mathcal{Y}(z)\right)^{\sigma_{3}/2}e^{2\phi_{{\overline{A}}}(z)^{1/2}\sigma_{3}}e^{i\pi\sigma_{3}/4}J_{\Psi}\left(\varphi_{{\overline{A}}}(z)\right)e^{-i\pi\sigma_{3}/4}e^{-2\varphi_{{\overline{A}}}(z)^{1/2}\sigma_{3}}\left(e^{2\widetilde{y}(z)}\mathcal{Y}(z)\right)^{\sigma_{3}/2}.$
From the properties of $\Psi$, for $z\in\omega_{1}\cup\omega_{2}$, we have:
$e^{i\pi\sigma_{3}/4}J_{\Psi}\left(\varphi_{{\overline{A}}}(z)\right)e^{-i\pi\sigma_{3}/4}=\begin{pmatrix}1&0\\ -i&1\end{pmatrix}.$
Inserting this into the equation for $J_{E_{{\overline{A}}}}(z)$, a straightforward calculation yields the desired jump. ∎
Remark 6.2.
We list here some properties of the map $\varphi_{{\overline{A}}}(z)^{1/2}$. The map $z\mapsto\varphi_{{\overline{A}}}(z)^{1/2}$ is a locally analytic map from $\mathcal{U}_{{\overline{A}}}$ to a neighborhood of 00 such that $\varphi_{{\overline{A}}}\left(\eta_{4}\cap\mathcal{U}_{{\overline{A}}}\right)\subset\mathbb{R}_{-}$. It is immediate to see that ${\overline{A}}$ is mapped to 00 under this map. Moreover, for $z\in\overline{\Omega_{2}^{+}}\cap\mathcal{U}_{{\overline{A}}}$, we have that $\varphi_{{\overline{A}}}(z)^{1/2}=-\ln(N){\widetilde{u}}(z)/4$. From lemma 4.1 it follows that ${\widetilde{u}}(z)$ is purely imaginary on the boundary of $\Omega_{2}^{+}$. This shows that $\varphi_{{\overline{A}}}(z)\in\mathbb{R}_{-}$ along $\eta_{4}\cap\mathcal{U}_{{\overline{A}}}$. Lastly, we choose $\omega_{1}$ and $\omega_{2}$ (within $\mathcal{U}_{{\overline{A}}}$) as the preimages of the rays with arguments ±$\pm 2\pi/3$, respectively.
@@ 6.3 Local parametrix near $z=A$
Recall that $\mathcal{U}_{A}$ is a small neighborhood of $z=A$, and we denote by $E_{A}$ the solution to the local problem near $z=A$. Inside $\mathcal{U}_{A}$, $E_{A}$ satisfies the jump relation
$E_{A+}(z)=E_{A+}(z)J_{E_{A}}(z),\quad J_{E_{A}}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{3}\\ \\[2.0pt] \mathfrak{M}^{-}(z),\quad&z\in\omega_{3}\cup\omega_{4}.\end{cases}$ (6.22)
@@
$E_{A}(z)=-Q_{A}(z)\sigma_{2}e^{i\pi\sigma_{3}/4}\Psi\big{(}\varphi_{A}(z)\big{)}e^{-i\pi\sigma_{3}/4}\sigma_{2}e^{-2\varphi_{A}(z)^{1/2}\sigma_{3}}\left(\frac{e^{2\widetilde{y}(z)}}{\mathcal{Y}(z)e^{4i\theta(z)}}\right)^{-\sigma_{3}/2}\,,$ (6.23)
satisfies the jump relation (6.22). Here, $Q_{A}$ is an analytic function inside $\mathcal{U}_{A}$ (to be determined later), and the map $\varphi_{A}$ is defined as
$\varphi_{A}(z)^{1/2}=-\ln(N)\frac{c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{S(s)}\,.$ (6.24)
@@
First, define
$\Psi_{A}(\xi)=-\sigma_{2}e^{i\pi\sigma_{3}/4}\Psi(\xi)e^{-i\pi\sigma_{3}/4}\sigma_{2}.$ (6.25)
The reader may verify that $\Psi_{A}$ satisfies the jump relation
$\left(\Psi_{A}\right)_{+}(\xi)=\left(\Psi_{A}\right)_{-}(\xi)\begin{cases}\begin{pmatrix}0&-i\\ -i&0\end{pmatrix},\quad&\xi\in(-\infty,0)\\ \\[4.0pt] \begin{pmatrix}1&-i\\ 0&1\end{pmatrix},\quad&\mbox{Arg}(\xi)=\pm 2\pi/3.\end{cases}$ (6.26)
The composition $\Psi_{A}(\varphi_{A}(z))$ possesses jumps across the contours $\eta_{3}$, $\omega_{3}$ and $\omega_{4}$. However, these contours are oriented in the opposite way as their images in the $\xi$ plane. The reader may verify that the jumps in the $z$ plane are the inverses of those in the $\xi$ plane:
$\left(\Psi_{A}(\varphi_{A}(z))\right)_{+}=\left(\Psi_{A}(\varphi_{A}(z))\right)_{-}\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{3}\\ \\[4.0pt] \begin{pmatrix}1&i\\ 0&1\end{pmatrix},\quad&z\in\omega_{3}\cup\omega_{4}.\end{cases}$ (6.27)
For example, as $z$ approaches $\hat{z}$ from the $+$ side of $\eta_{3}$, $\xi=\varphi_{A}(z)$ approaches $\varphi_{A}(\hat{z})$ from below, i.e. from the -- side of $\mathbb{R}_{-}$. Similarly, if $z$ approaches $\hat{z}$ from the -- side of $\eta_{3}$, then $\xi=\varphi_{A}(z)$ approaches $\varphi_{A}(\hat{z})$ from the $+$ side of $\mathbb{R}_{-}$. This means that
$\lim_{\begin{subarray}{c}z\to\hat{z}\\ z\in+\mbox{side of }\eta_{3}\end{subarray}}\Psi_{A}(\varphi_{A}(z))=\left(\Psi_{A}\right)_{-}(\varphi_{A}(\hat{z}))=\left(\Psi_{A}\right)_{+}(\varphi_{A}(\hat{z}))\left(\begin{array}[]{cc}0&i\\ i&0\end{array}\right)$ (6.30)
$=\lim_{\begin{subarray}{c}z\to\hat{z}\\ z\in-\mbox{side of }\eta_{3}\end{subarray}}\Psi_{A}(\varphi_{A}(z))\left(\begin{array}[]{cc}0&i\\ i&0\end{array}\right),$ (6.33)
which explains (6.27).
Now the definition (6.23) can be rewritten as
$E_{A}(z)=Q_{A}(z)\Psi_{A}(\varphi_{A}(z))e^{-2\varphi_{A}(z)^{1/2}\sigma_{3}}\left(\frac{e^{2\widetilde{y}(z)}}{\mathcal{Y}(z)e^{4i\theta(z)}}\right)^{-\sigma_{3}/2},$
and the known jump relations satisfied by the last two factors appearing on the right hand side can be used to verify that $E_{A}(z)$ satisfies the jump relations in (6.22). ∎